OpenStudy (anonymous):

The equation of line k is y = -1/5x – 1/8. What is the equation of the line that is perpendicular to line k and passes through (0, 0)?

OpenStudy (anonymous):

y = 5x Exp. The slope of given line is -(1/5) Slope of line perp. to it will be 5(because product of slopes of 2 perp. lines is -1) Equation of line passing through a given point is: y - a = m(x-b) (where a and b are the ordinate and abscissa) there fore the solution willl be y - 0 = 5(x-0) y = 5x